Circular No. 8664

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2000 C1 = P/2006 A3 (HERGENROTHER) E. J. Christensen reports the recovery of comet P/2000 C1 on 90-s

CCD exposures obtained on Jan. 6, 7, and 27 (initial observation tabulated below) with the Mt. Lemmon 1.5-m reflector. The Jan. 6 and 7 images were taken in very poor seeing and show no obvious sign of cometary activity. The Jan. 27 data were obtained in good seeing, but still show no indication of cometary activity. The object was consistently detected ∼ 3′ east of the predicted position for P/2000 C1; the indicated correction to the predictions on MPC 48384 and in the 2006 Comet Handbook is ΔT = −0.24 day.

2006 UT  α_{2000}  δ_{2000}  Mag.
Jan. 6.22144  5^{h}13^{m}36.^{s}01  +17^{o}22′15′′.6  21.4

The available astrometry, the following orbital elements (linking 2000 with 2006), and an ephemeris appear on MPEC 2006-B68.

Epoch = 2006 Nov. 1.0 TT

T = 2006 Nov. 6.6828 TT  ω = 51.^{o}2913
e = 0.407709  Ω = 127.0062 }2000.0
q = 2.088241 AU  i = 6.1079

a = 3.525702 AU n^{o} = 0.1488797 P = 6.620 years

SUPERNOVAE 2006Q, 2006R, 2006S Three additional apparent supernovae have been reported: 2006Q by

W. Li (LOSS/KAIT; cf. IAUC 8662) and the others by T. Puckett (cf. IAUC 8661; 2006R with A. Langoussis, and 2006S with R. Gagliano).

SN  2006 UT  α_{2000}  δ_{2000}  Mag.  Offset
2006Q Jan. 24.11 1^{h}38^{m}18.^{s}74 +35^{o}21′54′′.5 17.2 2′′.5 E, 1′′.0 N
2006R Jan. 26.47 16 23 20.82 +37 15 49.8 17.5 3′′.3 W, 19′′.3 N
2006S Jan. 26.44 12 45 39.10 +35 05 12.4 17.0 4′′.7 E, 10′′.7 N

Additional KAIT magnitudes of 2006Q in NGC 634: Jan. 13.11 UT, [19.5; 25.13, 17.2. SN 2006Q is probably a young type-II supernova (cf. CBET 381). Additional unfiltered CCD magnitudes of 2006R in NGC 6142: 2005 Oct. 15, [19.6; Jan. 27.36, 17.5. Additional approximate CCD magnitudes of 2006S in UGC 7934: 2005 Dec. 23, [19.8; 2006 Jan. 27.24, 17.0.

2006 January 27  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green